OpenStudy (anonymous):

if m angle1=(4x+12), angle2= (7x+3), and m angle 3= 6y then x =? m angle 1=?, and m angle 2 =? and y =?? help!!! plz

OpenStudy (anonymous):



OpenStudy (anonymous):

from figure we see angle 4 =180-angle 1 and by vertical angle theorem angle 2 and angle 4 are equal so 7x+3=180-4x-12 so x=15 and 4x+12=6y ---by vertical angle theorem so y=12 you can switch these values into angle 1 angl2 and get your answer

OpenStudy (anonymous):

angle 1=72 angle 2=108

OpenStudy (anonymous):

thank you!